OpenStudy (anonymous):

@Mathmatics What is the prime factorization of 216?

OpenStudy (anonymous):

\( 216=6^3 = (2\times3)^3 = 2^3 \times 3^3 \)

OpenStudy (anonymous):

Now how do you write this out? to see it work, like 216 = 6x2 = 12 x 6x2 = 6 x 12 Oh no I can not figure this one out?

OpenStudy (phi):

2*2*2*3*3*3

OpenStudy (anonymous):

Start with the smallest prime, work from there. When you get to a number that can't be divided by that, try the next. 216 / 2 = 108 / 2 = 54 / 2 = 27 / 3 = 9 / 3 = 3 / 3 = 1. So \(2\times 2 \times 2 \times 3 \times 3 \times 3 = {2^3}\times {3^3} = 216\)

OpenStudy (anonymous):

2x2=4 4x2=8 8x3=24 24x3=72 72x3=216

OpenStudy (anonymous):